$SPAD/src/input hyperbolicrules.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 Definition of Hyperbolic Functions

1.1 8.1 Hyperbolic sine of x

sinh(x) == \frac{e^{x} − e^{−x}}{2}

— * —

)set break resume
)spool hyperbolicrules.output
)set message test on
)set message auto off
)clear all
--S 1 of 298
sinhdef:=rule(sinh(x) == (e^x-e^(-x))/2)
--R
--R                    x    - x
--R                   e  - e
--R   (1)  sinh(x) == ---------
--R                       2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 1
--S 2 of 298
t1:=sinh(x) - (e^x-e^(-x))/2
--R
--R           x    - x
--R        - e  + e    + 2sinh(x)
--R   (2)  ----------------------
--R                   2
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 298
t2:=sinhdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 3

———-

1.2 8.2 Hyperbolic cosine of x

cosh(x) == \frac{e^{x} + e^{−x}}{2}

— * —

)clear all
--S 4 of 298
coshdef:=rule(cosh(x) == (e^x+e^(-x))/2)
--R
--R                    x    - x
--R                   e  + e
--R   (1)  cosh(x) == ---------
--R                       2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 4
--S 5 of 298
t1:=cosh(x) - (e^x+e^(-x))/2
--R
--R           x    - x
--R        - e  - e    + 2cosh(x)
--R   (2)  ----------------------
--R                   2
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 298
t2:=coshdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 6

———-

1.3 8.3 Hyperbolic tangent of x

tanh(x) == \frac{e^{x} − e^{−x}}{e^{x} + e^{−x}}

— * —

)clear all
--S 7 of 298
tanhdef:=rule(tanh(x) == (e^x-e*(-x))/(e^x+e*(-x)))
--R
--R                    x
--R                   e  + e x

--R   (1)  tanh(x) == --------
--R                    x
--R                   e  - e x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 7
--S 8 of 298
t1:=tanh(x) - (e^x-e*(-x))/(e^x+e*(-x))
--R
--R                      x
--R        (tanh(x) - 1)e  - e x tanh(x) - e x
--R   (2)  -----------------------------------
--R                       x
--R                      e  - e x
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 298
t2:=tanhdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 9

———-

1.4 8.4 Hyperbolic cotangent of x

coth(x) == \frac{e^{x} + e^{−x}}{e^{x} − e^{−x}}

— * —

)clear all
--S 10 of 298
cothdef:=rule(coth(x) == (e^x+e*(-x))/(e^x-e*(-x)))
--R
--R                    x
--R                   e  - e x
--R   (1)  coth(x) == --------
--R                    x
--R                   e  + e x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 10
--S 11 of 298
t1:=coth(x) - (e^x+e*(-x))/(e^x-e*(-x))

--R
--R                      x
--R        (coth(x) - 1)e  + e x coth(x) + e x
--R   (2)  -----------------------------------
--R                       x
--R                      e  + e x
--R                                                    Type: Expression(Integer)
--E 11
--S 12 of 298
t2:=cothdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 12

———-

1.5 8.5 Hyperbolic secant of x

sech(x) == \frac{2}{e^{x} + e^{−x}}

— * —

)clear all
--S 13 of 298
sechdef:=rule(sech(x) == 2/(e^x+e*(-x)))
--R
--R                       2
--R   (1)  sech(x) == --------
--R                    x
--R                   e  - e x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 13
--S 14 of 298
t1:=sech(x) - 2/(e^x+e*(-x))
--R
--R                x
--R        sech(x)e  - e x sech(x) - 2
--R   (2)  ---------------------------
--R                   x
--R                  e  - e x
--R                                                    Type: Expression(Integer)
--E 14

--S 15 of 298
t2:=sechdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 15

———-

1.6 8.6 Hyperbolic cosecant of x

csch(x) == \frac{2}{e^{x} − e^{−x}}

— * —

)clear all
--S 16 of 298
cschdef:=rule(csch(x) == 2/(e^x-e*(-x)))
--R
--R                       2
--R   (1)  csch(x) == --------
--R                    x
--R                   e  + e x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 16
--S 17 of 298
t1:=csch(x) - 2/(e^x-e*(-x))
--R
--R                x
--R        csch(x)e  + e x csch(x) - 2
--R   (2)  ---------------------------
--R                   x
--R                  e  + e x
--R                                                    Type: Expression(Integer)
--E 17
--S 18 of 298
t2:=cschdef t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 18

———-

2 Relationship among Hyperbolic Functions

2.1 8.7 tanh(x)

tanh(x) == \frac{sinh(x)}{cosh(x)}

— * —

)clear all
--S 19 of 298
tanhrule:=rule(tanh(x) == sinh(x)/cosh(x))
--R
--R                   sinh(x)
--R   (1)  tanh(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 19
--S 20 of 298
t1:=tanh(x) - sinh(x)/cosh(x)
--R
--R        cosh(x)tanh(x) - sinh(x)
--R   (2)  ------------------------
--R                 cosh(x)
--R                                                    Type: Expression(Integer)
--E 20
--S 21 of 298
t2:=tanhrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 21

———-

2.2 8.8 coth(x)

coth(x) == \frac{1}{tanh(x)}

== \frac{cosh(x)}{sinh(x)}

— * —

)clear all
--S 22 of 298
cothrule:=rule(coth(x) == 1/tanh(x))
--R
--R                      1
--R   (1)  coth(x) == -------
--R                   tanh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 22
--S 23 of 298
t1:=coth(x) - 1/tanh(x)
--R
--R        coth(x)tanh(x) - 1
--R   (2)  ------------------
--R              tanh(x)
--R                                                    Type: Expression(Integer)
--E 23
--S 24 of 298
t2:=cothrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 24
--S 25 of 298
cothrule2:=rule(coth(x) == cosh(x)/sinh(x))
--R
--R                   cosh(x)
--R   (4)  coth(x) == -------
--R                   sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 25
--S 26 of 298
t3:=coth(x) - cosh(x)/sinh(x)
--R
--R        coth(x)sinh(x) - cosh(x)
--R   (5)  ------------------------
--R                 sinh(x)
--R                                                    Type: Expression(Integer)
--E 26
--S 27 of 298
t4:=cothrule2 t3
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)

--E 27

———-

2.3 8.9 sech(x)

sech(x) == \frac{1}{cosh(x)}

— * —

)clear all
--S 28 of 298
sechrule:=rule(sech(x) == 1/cosh(x))
--R
--R                      1
--R   (1)  sech(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 28
--S 29 of 298
t1:=sech(x) - 1/cosh(x)
--R
--R        cosh(x)sech(x) - 1
--R   (2)  ------------------
--R              cosh(x)
--R                                                    Type: Expression(Integer)
--E 29
--S 30 of 298
t2:=sechrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 30

———-

2.4 8.10 csch(x)

csch(x) == \frac{1}{sinh(x)}

— * —

--S 31 of 298
cschrule:=rule(csch(x) == 1/sinh(x))
--R
--R                      1
--R   (1)  csch(x) == -------
--R                   sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 31
--S 32 of 298
t1:=csch(x) - 1/sinh(x)
--R
--R        csch(x)sinh(x) - 1
--R   (2)  ------------------
--R              sinh(x)
--R                                                    Type: Expression(Integer)
--E 32
--S 33 of 298
t2:=cschrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 33

———-

2.5 8.11 cosh^{2}(x) − sinh^{2}(x)

cosh^{2}(x) − sinh^{2}(x) == 1

— * —

)clear all
--S 34 of 298
coshsinh:=rule(cosh(x)^2-sinh(x)^2 == 1)
--R
--R                 2          2
--I   (1)  - sinh(x)  + cosh(x)  + %Y == %Y + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 34
--S 35 of 298
t1:=cosh(x)^2-sinh(x)^2 - 1

--R
--R                 2          2
--R   (2)  - sinh(x)  + cosh(x)  - 1
--R                                                    Type: Expression(Integer)
--E 35
--S 36 of 298
t2:=coshsinh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 36

———-

2.6 8.12 sech^{2}(x) + tanh^{2}(x)

sech^{2}(x) + tanh^{2}(x) == 1

— * —

)clear all
--S 37 of 298
sechtanh:=rule(sech(x)^2+tanh(x)^2 == 1)
--R
--R               2          2
--I   (1)  tanh(x)  + sech(x)  + %Z == %Z + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 37
--S 38 of 298
t1:=sech(x)^2+tanh(x)^2 - 1
--R
--R               2          2
--R   (2)  tanh(x)  + sech(x)  - 1
--R                                                    Type: Expression(Integer)
--E 38
--S 39 of 298
t2:=sechtanh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 39

———-

2.7 8.13 coth^{2}(x) − csch^{2}(x)

coth^{2}(x) − csch^{2}(x) == 1

— * —

)clear all
--S 40 of 298
cothcsch:=rule(coth(x)^2-csch(x)^2 == 1)
--R
--R                 2          2
--I   (1)  - csch(x)  + coth(x)  + %BA == %BA + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 40
--S 41 of 298
t1:=coth(x)^2-csch(x)^2 - 1
--R
--R                 2          2
--R   (2)  - csch(x)  + coth(x)  - 1
--R                                                    Type: Expression(Integer)
--E 41
--S 42 of 298
t2:=cothcsch t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 42

———-

3 Functions of Negative Arguments

Axiom already knows these simplifications

3.1 8.14 sinh(−x)

sinh(−x) == − sinh(x)

— * —

--S 43 of 298
sinh(-x)
--R
--R   (1)  - sinh(x)
--R                                                    Type: Expression(Integer)
--E 43

———-

3.2 8.15 cosh(−x)

cosh(−x) == cosh(x)

— * —

)clear all
--S 44 of 298
cosh(-x)
--R
--R   (1)  cosh(x)
--R                                                    Type: Expression(Integer)
--E 44

———-

3.3 8.16 tanh(−x)

tanh(−x) == − tanh(x)

— * —

)clear all
--S 45 of 298
tanh(-x)
--R
--R   (1)  - tanh(x)
--R                                                    Type: Expression(Integer)
--E 45

———-

3.4 8.17 csch(−x)

csch(−x) == − csch(x)

— * —

)clear all
--S 46 of 298
csch(-x)
--R
--R   (1)  - csch(x)
--R                                                    Type: Expression(Integer)
--E 46

———-

3.5 8.18 sech(−x)

sech(−x) == sech(x)

— * —

)clear all
--S 47 of 298
sech(-x)
--R
--R   (1)  sech(x)
--R                                                    Type: Expression(Integer)
--E 47

———-

3.6 8.19 coth(−x)

coth(−x) == − coth(x)

— * —

--S 48 of 298
coth(-x)
--R
--R   (1)  - coth(x)
--R                                                    Type: Expression(Integer)
--E 48

———-

4 Addition Formulas

4.1 8.20 sinh(x ± y)

sinh(x ± y) == sinh(x) cosh(y) ± cosh(x) sinh(y)

— * —

)clear all
--S 49 of 298
sinhadd:=rule(sinh(x+y) == sinh(x)*cosh(y)+cosh(x)*sinh(y))
--R
--R   (1)  sinh(y + x) == cosh(x)sinh(y) + cosh(y)sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 49
--S 50 of 298
t1:=sinh(x+y) - (sinh(x)*cosh(y)+cosh(x)*sinh(y))
--R
--R   (2)  sinh(y + x) - cosh(x)sinh(y) - cosh(y)sinh(x)
--R                                                    Type: Expression(Integer)
--E 50
--S 51 of 298
t2:=sinhadd t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 51
--S 52 of 298
sinhsub:=rule(sinh(x-y) == sinh(x)*cosh(y)-cosh(x)*sinh(y))
--R
--I   (4)  - %T sinh(y - x) == - %T cosh(x)sinh(y) + %T cosh(y)sinh(x)

--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 52
--S 53 of 298
t3:=sinh(x-y) - (sinh(x)*cosh(y)-cosh(x)*sinh(y))
--R
--R   (5)  cosh(x)sinh(y) - sinh(y - x) - cosh(y)sinh(x)
--R                                                    Type: Expression(Integer)
--E 53
--S 54 of 298
t4:=sinhsub t3
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E 54

———-

4.2 8.21 cosh(x ± y)

cosh(x ± y) == cosh(x) cosh(y) ± sinh(x) sinh(y)

— * —

)clear all
--S 55 of 298
coshadd:=rule(cosh(x+y) == cosh(x)*cosh(y)+sinh(x)*sinh(y))
--R
--R   (1)  cosh(y + x) == sinh(x)sinh(y) + cosh(x)cosh(y)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 55
--S 56 of 298
t1:=cosh(x+y) - (cosh(x)*cosh(y)+sinh(x)*sinh(y))
--R
--R   (2)  - sinh(x)sinh(y) + cosh(y + x) - cosh(x)cosh(y)
--R                                                    Type: Expression(Integer)
--E 56
--S 57 of 298
t2:=coshadd t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 57

--S 58 of 298
coshsub:=rule(cosh(x-y) == cosh(x)*cosh(y)-sinh(x)*sinh(y))
--R
--R   (4)  cosh(y - x) == - sinh(x)sinh(y) + cosh(x)cosh(y)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 58
--S 59 of 298
t3:=cosh(x-y) - (cosh(x)*cosh(y)-sinh(x)*sinh(y))
--R
--R   (5)  sinh(x)sinh(y) - cosh(x)cosh(y) + cosh(y - x)
--R                                                    Type: Expression(Integer)
--E 59
--S 60 of 298
t4:=coshsub t3
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E 60

———-

4.3 8.22 tanh(x ± y)

tanh(x ± y) == \frac{tanh(x) ± tanh(y)}{1 ± tanh(x) tanh(y)}

— * —

)clear all
--S 61 of 298
tanhadd:=rule(tanh(x+y) == (tanh(x)+tanh(y))/(1+tanh(x)*tanh(y)))
--R
--R                        tanh(y) + tanh(x)
--R   (1)  tanh(y + x) == ------------------
--R                       tanh(x)tanh(y) + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 61
--S 62 of 298
t1:=tanh(x+y) - (tanh(x)+tanh(y))/(1+tanh(x)*tanh(y))
--R
--R        (tanh(x)tanh(y) + 1)tanh(y + x) - tanh(y) - tanh(x)
--R   (2)  ---------------------------------------------------
--R                         tanh(x)tanh(y) + 1

--R                                                    Type: Expression(Integer)
--E 62
--S 63 of 298
t2:=tanhadd t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 63
--S 64 of 298
tanhneg:=rule(tanh(x-y) == (tanh(x)-tanh(y))/(1-tanh(x)*tanh(y)))
--R
--I                            %V tanh(y) - %V tanh(x)
--I   (4)  - %V tanh(y - x) == -----------------------
--R                               tanh(x)tanh(y) - 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 64
--S 65 of 298
t3:=tanh(x-y) - (tanh(x)-tanh(y))/(1-tanh(x)*tanh(y))
--R
--R        (- tanh(x)tanh(y - x) - 1)tanh(y) + tanh(y - x) + tanh(x)
--R   (5)  ---------------------------------------------------------
--R                            tanh(x)tanh(y) - 1
--R                                                    Type: Expression(Integer)
--E 65

———-

This loops forever.

— * —

--S 66 of 298
-- t4:=tanhneg t3
--E 66

———-

4.4 8.23 coth(x ± y)

coth(x ± y) == \frac{coth(x) coth(y) ± 1}{coth(y) ± coth(x)}

— * —

)clear all

--S 67 of 298
cothadd:=rule(coth(x+y) == (coth(x)*coth(y)+1)/(coth(y)+coth(x)))
--R
--R                       coth(x)coth(y) + 1
--R   (1)  coth(y + x) == ------------------
--R                        coth(y) + coth(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 67
--S 68 of 298
t1:=coth(x+y) - (coth(x)*coth(y)+1)/(coth(y)+coth(x))
--R
--R        (coth(y) + coth(x))coth(y + x) - coth(x)coth(y) - 1
--R   (2)  ---------------------------------------------------
--R                         coth(y) + coth(x)
--R                                                    Type: Expression(Integer)
--E 68
--S 69 of 298
t2:=cothadd t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 69
--S 70 of 298
cothneg:=rule(coth(x-y) == (coth(x)*coth(y)-1)/(coth(y)-coth(x)))
--R
--I                            %W coth(x)coth(y) - %W
--I   (4)  - %W coth(y - x) == ----------------------
--R                               coth(y) - coth(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 70
--S 71 of 298
t3:=coth(x-y) - (coth(x)*coth(y)-1)/(coth(y)-coth(x))
--R
--R        (- coth(y - x) - coth(x))coth(y) + coth(x)coth(y - x) + 1
--R   (5)  ---------------------------------------------------------
--R                            coth(y) - coth(x)
--R                                                    Type: Expression(Integer)
--E 71

———-

This loops forever — * —

--S 72 of 298
--t4:=cothneg t3

--E 72

———-

5 Double Angle Formulas

5.1 8.24 sinh(2x)

sinh(2x) == 2 sinh(x) cosh(x)

— * —

)clear all
--S 73 of 298
sinh2x:=rule(sinh(2*x) == 2*sinh(x)*cosh(x))
--R
--R   (1)  sinh(2x) == 2cosh(x)sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 73
--S 74 of 298
t1:=sinh(2*x) - 2*sinh(x)*cosh(x)
--R
--R   (2)  sinh(2x) - 2cosh(x)sinh(x)
--R                                                    Type: Expression(Integer)
--E 74
--S 75 of 298
t2:=sinh2x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 75

———-

5.2 8.25 cosh(2x)

cosh(2x) == cosh^{2}(x) + sinh^{2}(x)
== 2 cosh^{2}(x) − 1
== 1 + 2 sinh^{2}(x)

— * —

--S 76 of 298
cosh2x:=rule(cosh(2*x) == cosh(x)^2+sinh(x)^2)
--R
--R                           2          2
--R   (1)  cosh(2x) == sinh(x)  + cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 76
--S 77 of 298
t1:=cosh(2*x) - (cosh(x)^2+sinh(x)^2)
--R
--R                 2                     2
--R   (2)  - sinh(x)  + cosh(2x) - cosh(x)
--R                                                    Type: Expression(Integer)
--E 77
--S 78 of 298
t2:=cosh2x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 78
--S 79 of 298
cosh2x2:=rule(cosh(2*x) == 2*cosh(x)^2-1)
--R
--R                            2
--R   (4)  cosh(2x) == 2cosh(x)  - 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 79
--S 80 of 298
t3:=cosh(2*x) - (2*cosh(x)^2-1)
--R
--R                           2
--R   (5)  cosh(2x) - 2cosh(x)  + 1
--R                                                    Type: Expression(Integer)
--E 80
--S 81 of 298
t4:=cosh2x2 t3
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E 81
--S 82 of 298
cosh2x3:=rule(cosh(2*x) == 1+2*sinh(x)^2)

--R
--R                            2
--R   (7)  cosh(2x) == 2sinh(x)  + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 82
--S 83 of 298
t5:=cosh(2*x) - (1+2*sinh(x)^2)
--R
--R                  2
--R   (8)  - 2sinh(x)  + cosh(2x) - 1
--R                                                    Type: Expression(Integer)
--E 83
--S 84 of 298
t6:=cosh2x3 t5
--R
--R   (9)  0
--R                                                    Type: Expression(Integer)
--E 84

———-

5.3 8.26 tanh(2x)

tanh(2x) == \frac{2 tanh(x)}{1 + tanh^{2}(x)}

— * —

)clear all
--S 85 of 298
tanh2x:=rule(tanh(2*x) == (2*tanh(x))/(1+tanh(x)^2))
--R
--R                      2tanh(x)
--R   (1)  tanh(2x) == ------------
--R                           2
--R                    tanh(x)  + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 85
--S 86 of 298
t1:=tanh(2*x) - (2*tanh(x))/(1+tanh(x)^2)
--R
--R                2
--R        (tanh(x)  + 1)tanh(2x) - 2tanh(x)
--R   (2)  ---------------------------------

--R                          2
--R                   tanh(x)  + 1
--R                                                    Type: Expression(Integer)
--E 86
--S 87 of 298
t2:=tanh2x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 87

———-

6 Half Angle Formulas

6.1 8.27 sinh (\frac{x}{2})

sinh (\frac{x}{2}) == ±\sqrt{\frac{cosh(x) − 1}{2}} [+ if x > 0, −if x < 0]

if x > 0 then — * —

)clear all
--S 88 of 298
sinhhalf:=rule(sinh(x/2) == sqrt((cosh(x)-1)/2))
--R
--R                    +-----------+
--R             x     \|cosh(x) - 1
--R   (1)  sinh(-) == --------------
--R             2           +-+
--R                        \|2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 88
--S 89 of 298
t1:=sinh(x/2) - sqrt((cosh(x)-1)/2)
--R
--R           +-----------+    +-+     x
--R        - \|cosh(x) - 1  + \|2 sinh(-)
--R                                    2
--R   (2)  ------------------------------
--R                      +-+
--R                     \|2
--R                                                    Type: Expression(Integer)

--E 89
--S 90 of 298
t2:=sinhhalf t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 90

———-

if x < 0 then — * —

)clear all
--S 91 of 298
sinhhalfneg:=rule(sinh(x/2) == -sqrt((cosh(x)-1)/2))
--R
--R                      +-----------+
--R             x       \|cosh(x) - 1
--R   (1)  sinh(-) == - --------------
--R             2             +-+
--R                          \|2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 91
--S 92 of 298
t1:=sinh(x/2) - -sqrt((cosh(x)-1)/2)
--R
--R         +-----------+    +-+     x
--R        \|cosh(x) - 1  + \|2 sinh(-)
--R                                  2
--R   (2)  ----------------------------
--R                     +-+
--R                    \|2
--R                                                    Type: Expression(Integer)
--E 92
--S 93 of 298
t2:=sinhhalfneg t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 93

———-

6.2 8.28 cosh (\frac{x}{2})

cosh (\frac{x}{2}) == \sqrt{\frac{cosh(x) + 1}{2}}

— * —

)clear all
--S 94 of 298
coshhalf:=rule(cosh(x/2) == sqrt((cosh(x)+1)/2))
--R
--R                    +-----------+
--R             x     \|cosh(x) + 1
--R   (1)  cosh(-) == --------------
--R             2           +-+
--R                        \|2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 94
--S 95 of 298
t1:=cosh(x/2) - sqrt((cosh(x)+1)/2)
--R
--R           +-----------+    +-+     x
--R        - \|cosh(x) + 1  + \|2 cosh(-)
--R                                    2
--R   (2)  ------------------------------
--R                      +-+
--R                     \|2
--R                                                    Type: Expression(Integer)
--E 95
--S 96 of 298
t2:=coshhalf t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 96

———-

6.3 8.29 tanh (\frac{x}{2})

tanh (\frac{x}{2}) == ±\sqrt{\frac{cosh(x) − 1}{cosh(x) + 1}} [+ if x > 0, −if x < 0]

== \frac{sinh(x)}{cosh(x) + 1}

== \frac{cosh(x) − 1}{sinh(x)}

if x > 0 — * —

)clear all
--S 97 of 298
tanhhalf:=rule(tanh(x/2) == sqrt((cosh(x)-1)/(cosh(x)+1)))
--R
--R                    +-----------+
--R             x      |cosh(x) - 1
--R   (1)  tanh(-) ==  |-----------
--R             2     \|cosh(x) + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 97
--S 98 of 298
t1:=tanh(x/2) -sqrt((cosh(x)-1)/(cosh(x)+1))
--R
--R           +-----------+
--R           |cosh(x) - 1         x
--R   (2)  -  |-----------  + tanh(-)
--R          \|cosh(x) + 1         2
--R                                                    Type: Expression(Integer)
--E 98
--S 99 of 298
t2:=tanhhalf t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 99

———-

if x < 0 — * —

)clear all
--S 100 of 298
tanhhalfneg:=rule(tanh(x/2) == -sqrt((cosh(x)-1)/(cosh(x)+1)))
--R
--R                      +-----------+
--R             x        |cosh(x) - 1
--R   (1)  tanh(-) == -  |-----------

--R             2       \|cosh(x) + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 100
--S 101 of 298
t1:=tanh(x/2) - -sqrt((cosh(x)-1)/(cosh(x)+1))
--R
--R         +-----------+
--R         |cosh(x) - 1         x
--R   (2)   |-----------  + tanh(-)
--R        \|cosh(x) + 1         2
--R                                                    Type: Expression(Integer)
--E 101
--S 102 of 298
t2:=tanhhalfneg t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 102
)clear all
--S 103 of 298
tanhhalf2:=rule(tanh(x/2) == sinh(x)/(cosh(x)+1))
--R
--R             x       sinh(x)
--R   (1)  tanh(-) == -----------
--R             2     cosh(x) + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 103
--S 104 of 298
t1:=tanh(x/2) - sinh(x)/(cosh(x)+1)
--R
--R                          x
--R        (cosh(x) + 1)tanh(-) - sinh(x)
--R                          2
--R   (2)  ------------------------------
--R                  cosh(x) + 1
--R                                                    Type: Expression(Integer)
--E 104
--S 105 of 298
t2:=tanhhalf2 t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 105

--S 106 of 298
tanhhalf3:=rule(tanh(x/2) == (cosh(x)-1)/sinh(x))
--R
--R             x     cosh(x) - 1
--R   (1)  tanh(-) == -----------
--R             2       sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 106
--S 107 of 298
t1:=tanh(x/2) - (cosh(x)-1)/sinh(x)
--R
--R                    x
--R        sinh(x)tanh(-) - cosh(x) + 1
--R                    2
--R   (2)  ----------------------------
--R                   sinh(x)
--R                                                    Type: Expression(Integer)
--E 107
--S 108 of 298
t2:=tanhhalf3 t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 108

———-

7 Multiple Angle Formulas

7.1 8.30 sinh(3x)

sinh(3x) == 3 sinh(x) + 4 sinh^{3}(x)

— * —

)clear all
--S 109 of 298
sinh3x:=rule(sinh(3*x) == 3*sinh(x)+4*sinh(x)^3)
--R
--R                            3
--R   (1)  sinh(3x) == 4sinh(x)  + 3sinh(x)

--S 110 of 298
t1:=sinh(3*x) - (3*sinh(x)+4*sinh(x)^3)
--R
--R                           3
--R   (2)  sinh(3x) - 4sinh(x)  - 3sinh(x)
--R                                                    Type: Expression(Integer)
--E 110
--S 111 of 298
t2:=sinh3x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 111

———-

7.2 8.31 cosh(3x)

cosh(3x) == 4 cosh^{3}(x) − 3 cosh(x)

— * —

)clear all
--S 112 of 298
cosh3x:=rule(cosh(3*x) == 4*cosh(x)^3-3*cosh(x))
--R
--R                            3
--R   (1)  cosh(3x) == 4cosh(x)  - 3cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 112
--S 113 of 298
t1:=cosh(3*x) - (4*cosh(x)^3-3*cosh(x))
--R
--R                           3
--R   (2)  cosh(3x) - 4cosh(x)  + 3cosh(x)
--R                                                    Type: Expression(Integer)
--E 113
--S 114 of 298
t2:=cosh3x t1
--R

--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 114

———-

7.3 8.32 tanh(3x)

tanh(3x) == \frac{3 tanh(x) + tanh^{3}(x)}{1 + 3 tanh^{2}(x)}

— * —

)clear all
--S 115 of 298
tanh3x:=rule(tanh(3*x) == (3*tanh(x)+tanh(x)^3)/(1+3*tanh(x)^2))
--R
--R                           3
--R                    tanh(x)  + 3tanh(x)
--R   (1)  tanh(3x) == -------------------
--R                               2
--R                       3tanh(x)  + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 115
--S 116 of 298
t1:=tanh(3*x) - (3*tanh(x)+tanh(x)^3)/(1+3*tanh(x)^2)
--R
--R                 2                       3
--R        (3tanh(x)  + 1)tanh(3x) - tanh(x)  - 3tanh(x)
--R   (2)  ---------------------------------------------
--R                                2
--R                        3tanh(x)  + 1
--R                                                    Type: Expression(Integer)
--E 116
--S 117 of 298
t2:=tanh3x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 117

———-

7.4 8.33 sinh(4x)

sinh(4x) == 8 sinh^{3}(x) cosh(x) + 4 sinh(x) cosh(x)

— * —

)clear all
--S 118 of 298
sinh4x:=rule(sinh(4*x) == 8*sinh(x)^3*cosh(x)+4*sinh(x)*cosh(x))
--R
--R                                   3
--R   (1)  sinh(4x) == 8cosh(x)sinh(x)  + 4cosh(x)sinh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 118
--S 119 of 298
t1:=sinh(4*x) - (8*sinh(x)^3*cosh(x)+4*sinh(x)*cosh(x))
--R
--R                                  3
--R   (2)  sinh(4x) - 8cosh(x)sinh(x)  - 4cosh(x)sinh(x)
--R                                                    Type: Expression(Integer)
--E 119
--S 120 of 298
t2:=sinh4x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 120

———-

7.5 8.34 cosh(4x)

cosh(4x) == 8 cosh^{4}(x) − 8 cosh^{2}(x) + 1

— * —

)clear all
--S 121 of 298
cosh4x:=rule(cosh(4*x) == 8*cosh(x)^4-8*cosh(x)^2+1)
--R
--R                            4           2

--R   (1)  cosh(4x) == 8cosh(x)  - 8cosh(x)  + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 121
--S 122 of 298
t1:=cosh(4*x) - (8*cosh(x)^4-8*cosh(x)^2+1)
--R
--R                           4           2
--R   (2)  cosh(4x) - 8cosh(x)  + 8cosh(x)  - 1
--R                                                    Type: Expression(Integer)
--E 122
--S 123 of 298
t2:=cosh4x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 123

———-

7.6 8.35 tanh(4x)

tanh(4x) == \frac{4 tanh(x) + 4 tanh^{3}(x)}{1 + 6 tanh^{2}(x) + tanh^{4}(x)}

— * —

)clear all
--S 124 of 298
tanh4x:=rule(tanh(4*x) == (4*tanh(x)+4*tanh(x)^3)/(1+6*tanh(x)^2+tanh(x)^4))
--R
--R                              3
--R                      4tanh(x)  + 4tanh(x)
--R   (1)  tanh(4x) == ------------------------
--R                           4           2
--R                    tanh(x)  + 6tanh(x)  + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 124
--S 125 of 298
t1:=tanh(4*x) - (4*tanh(x)+4*tanh(x)^3)/(1+6*tanh(x)^2+tanh(x)^4)
--R
--R                4           2                        3
--R        (tanh(x)  + 6tanh(x)  + 1)tanh(4x) - 4tanh(x)  - 4tanh(x)
--R   (2)  ---------------------------------------------------------
--R                                4           2

--R                         tanh(x)  + 6tanh(x)  + 1
--R                                                    Type: Expression(Integer)
--E 125
--S 126 of 298
t2:=tanh4x t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 126

———-

8 Powers of Hyperbolic Functions

8.1 8.36 sinh^{2}(x)

sinh^{2}(x) == \frac{1}{2} cosh(2x) − \frac{1}{2}

— * —

)clear all
--S 127 of 298
sinhsquared:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (1)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 127
--S 128 of 298
t1:=sinh(x)^2 - (1/2*cosh(2*x)-1/2)
--R
--R                2
--R        2sinh(x)  - cosh(2x) + 1
--R   (2)  ------------------------
--R                    2
--R                                                    Type: Expression(Integer)
--E 128
--S 129 of 298
t2:=sinhsquared t1
--R
--R   (3)  0

———-

8.2 8.37 cosh^{2}(x)

cosh^{2}(x) == \frac{1}{2} cosh(2x) + \frac{1}{2}

— * —

)clear all
--S 130 of 298
coshsquared:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--E 130
--S 131 of 298
t1:=cosh(x)^2 - (1/2*cosh(2*x)+1/2)
--E 131
--S 132 of 298
t2:=coshsquared t1
--E 132

———-

8.3 8.38 sinh^{3}(x)

sinh^{3}(x) == \frac{1}{4} sinh(3x) − \frac{3}{4} sinh(x)

— * —

)clear all
--S 133 of 298
sinhcubed:=rule(sinh(x)^3 == 1/4*sinh(3*x)-3/4*sinh(x))
--R
--R               3    sinh(3x) - 3sinh(x)
--R   (1)  sinh(x)  == -------------------
--R                             4
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 133

--S 134 of 298
t1:=sinh(x)^3 - (1/4*sinh(3*x)-3/4*sinh(x))
--R
--R                             3
--R        - sinh(3x) + 4sinh(x)  + 3sinh(x)
--R   (2)  ---------------------------------
--R                        4
--R                                                    Type: Expression(Integer)
--E 134
--S 135 of 298
t2:=sinhcubed t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 135

———-

8.4 8.39 cosh^{3}(x)

cosh^{3}(x) == \frac{1}{4} cosh(3x) + \frac{3}{4} cosh(x)

— * —

)clear all
--S 136 of 298
coshcubed:=rule(cosh(x)^3 == 1/4*cosh(3*x)+3/4*cosh(x))
--R
--R               3    cosh(3x) + 3cosh(x)
--R   (1)  cosh(x)  == -------------------
--R                             4
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 136
--S 137 of 298
t1:=cosh(x)^3 - (1/4*cosh(3*x)+3/4*cosh(x))
--R
--R                             3
--R        - cosh(3x) + 4cosh(x)  - 3cosh(x)
--R   (2)  ---------------------------------
--R                        4
--R                                                    Type: Expression(Integer)
--E 137
--S 138 of 298

t2:=coshcubed t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 138

———-

8.5 8.40 sinh^{4}(x)

sinh^{4}(x) == \frac{3}{8} − \frac{1}{2} cosh(2x) + \frac{1}{8} cosh(4x)

— * —

)clear all
--S 139 of 298
sinhfourth:=rule(sinh(x)^4 == 3/8-1/2*cosh(2*x)+1/8*cosh(4*x))
--R
--R               4    cosh(4x) - 4cosh(2x) + 3
--R   (1)  sinh(x)  == ------------------------
--R                                8
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 139
--S 140 of 298
t1:=sinh(x)^4 - (3/8-1/2*cosh(2*x)+1/8*cosh(4*x))
--R
--R                4
--R        8sinh(x)  - cosh(4x) + 4cosh(2x) - 3
--R   (2)  ------------------------------------
--R                          8
--R                                                    Type: Expression(Integer)
--E 140
--S 141 of 298
t2:=sinhfourth t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 141

———-

8.6 8.41 cosh^{4}(x)

cosh^{4}(x) == \frac{3}{8} + \frac{1}{2} cosh(2x) + \frac{1}{8} cosh(4x)

— * —

)clear all
--S 142 of 298
coshfourth:=rule(cosh(x)^4 == 3/8+1/2*cosh(2*x)+1/8*cosh(4*x))
--R
--R               4    cosh(4x) + 4cosh(2x) + 3
--R   (1)  cosh(x)  == ------------------------
--R                                8
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 142
--S 143 of 298
t1:=cosh(x)^4 - (3/8+1/2*cosh(2*x)+1/8*cosh(4*x))
--R
--R                                         4
--R        - cosh(4x) - 4cosh(2x) + 8cosh(x)  - 3
--R   (2)  --------------------------------------
--R                           8
--R                                                    Type: Expression(Integer)
--E 143
--S 144 of 298
t2:=coshfourth t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 144

———-

9 Sum, Difference, and Product of Hyperbolic Products

9.1 8.42 sinh(x) + sinh(y)

sinh(x) + sinh(y) == 2 sinh(\frac{1}{2}(x + y)) cosh(\frac{1}{2}(x − y))

— * —

--S 145 of 298
sinhplussinh:=rule(sinh(x)+sinh(y) == 2*sinh(1/2*(x+y))*cosh(1/2*(x-y)))
--R
--R                                        y - x      y + x
--I   (1)  sinh(y) + sinh(x) + %M == 2cosh(-----)sinh(-----) + %M
--R                                          2          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 145
--S 146 of 298
t1:=sinh(x)+sinh(y) - 2*sinh(1/2*(x+y))*cosh(1/2*(x-y))
--R
--R                        y - x      y + x
--R   (2)  sinh(y) - 2cosh(-----)sinh(-----) + sinh(x)
--R                          2          2
--R                                                    Type: Expression(Integer)
--E 146
--S 147 of 298
t2:=sinhplussinh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 147

———-

9.2 8.43 sinh(x) − sinh(y)

sinh(x) − sinh(y) == 2 cosh(\frac{1}{2}(x + y)) sinh(\frac{1}{2}(x − y))

— * —

)clear all
--S 148 of 298
sinhminussinh:=rule(sinh(x)-sinh(y) == 2*cosh(1/2*(x+y))*sinh(1/2*(x-y)))
--R
--R                                            y + x      y - x
--I   (1)  - sinh(y) + sinh(x) + %N == - 2cosh(-----)sinh(-----) + %N
--R                                              2          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 148
--S 149 of 298

t1:=sinh(x)-sinh(y) - 2*cosh(1/2*(x+y))*sinh(1/2*(x-y))
--R
--R                          y + x      y - x
--R   (2)  - sinh(y) + 2cosh(-----)sinh(-----) + sinh(x)
--R                            2          2
--R                                                    Type: Expression(Integer)
--E 149
--S 150 of 298
t2:=sinhminussinh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 150

———-

9.3 8.44 cosh(x) + cosh(y)

cosh(x) + cosh(y) == 2 cosh(\frac{1}{2}(x + y)) cosh(\frac{1}{2}(x − y))

— * —

)clear all
--S 151 of 298
coshpluscosh:=rule(cosh(x)+cosh(y) == 2*cosh(1/2*(x+y))*cosh(1/2*(x-y)))
--R
--R                                        y - x      y + x
--I   (1)  cosh(y) + cosh(x) + %O == 2cosh(-----)cosh(-----) + %O
--R                                          2          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 151
--S 152 of 298
t1:=cosh(x)+cosh(y) - 2*cosh(1/2*(x+y))*cosh(1/2*(x-y))
--R
--R                        y - x      y + x
--R   (2)  cosh(y) - 2cosh(-----)cosh(-----) + cosh(x)
--R                          2          2
--R                                                    Type: Expression(Integer)
--E 152
--S 153 of 298
t2:=coshpluscosh t1
--R
--R   (3)  0

———-

9.4 8.45 cosh(x) − cosh(y)

cosh(x) − cosh(y) == 2 sinh(\frac{1}{2}(x + y)) sinh(\frac{1}{2}(x − y))

— * —

)clear all
--S 154 of 298
coshminuscosh:=rule(cosh(x)-cosh(y) == 2*sinh(1/2*(x+y))*sinh(1/2*(x-y)))
--R
--R                                            y - x      y + x
--I   (1)  - cosh(y) + cosh(x) + %P == - 2sinh(-----)sinh(-----) + %P
--R                                              2          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 154
--S 155 of 298
t1:=cosh(x)-cosh(y) - 2*sinh(1/2*(x+y))*sinh(1/2*(x-y))
--R
--R              y - x      y + x
--R   (2)  2sinh(-----)sinh(-----) - cosh(y) + cosh(x)
--R                2          2
--R                                                    Type: Expression(Integer)
--E 155
--S 156 of 298
t2:=coshminuscosh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 156

———-

9.5 8.46 sinh(x)sinh(y)

sinh(x)sinh(y) == \frac{1}{2}( cosh(x + y) − cosh(x − y))

— * —

--S 157 of 298
sinhtimessinh:=rule(sinh(x)*sinh(y) == 1/2*(cosh(x+y)-cosh(x-y)))
--R
--I                             %Q cosh(y + x) - %Q cosh(y - x)
--I   (1)  %Q sinh(x)sinh(y) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 157
--S 158 of 298
t1:=sinh(x)*sinh(y) - (1/2*(cosh(x+y)-cosh(x-y)))
--R
--R        2sinh(x)sinh(y) - cosh(y + x) + cosh(y - x)
--R   (2)  -------------------------------------------
--R                             2
--R                                                    Type: Expression(Integer)
--E 158
--S 159 of 298
t2:=sinhtimessinh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 159

———-

9.6 8.47 cosh(x)cosh(y)

cosh(x)cosh(y) == \frac{1}{2}( cosh(x + y) + cosh(x − y))

— * —

)clear all
--S 160 of 298
coshtimescosh:=rule(cosh(x)*cosh(y)==1/2*(cosh(x+y)+cosh(x-y)))
--R
--I                             %R cosh(y + x) + %R cosh(y - x)
--I   (1)  %R cosh(x)cosh(y) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 160
--S 161 of 298

t1:=cosh(x)*cosh(y) - 1/2*(cosh(x+y)+cosh(x-y))
--R
--R        - cosh(y + x) + 2cosh(x)cosh(y) - cosh(y - x)
--R   (2)  ---------------------------------------------
--R                              2
--R                                                    Type: Expression(Integer)
--E 161
--S 162 of 298
t2:=coshtimescosh t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 162

———-

9.7 8.48 sinh(x)cosh(y)

sinh(x)cosh(y) == \frac{1}{2}( sinh(x + y) + sinh(x − y))

— * —

)clear all
--S 163 of 298
sinhtimescosh:=rule(sinh(x)*cosh(y) == 1/2*(sinh(x+y)+sinh(x-y)))
--R
--I                             %S sinh(y + x) - %S sinh(y - x)
--I   (1)  %S cosh(y)sinh(x) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 163
--S 164 of 298
t1:=sinh(x)*cosh(y) - 1/2*(sinh(x+y)+sinh(x-y))
--R
--R        - sinh(y + x) + sinh(y - x) + 2cosh(y)sinh(x)
--R   (2)  ---------------------------------------------
--R                              2
--R                                                    Type: Expression(Integer)
--E 164
--S 165 of 298
t2:=sinhtimescosh t1
--R
--R   (3)  0

———-

10 Inverse Hyperbolic Functions

10.1 8.55 sinh^{−1}(x)

asinh(x) == log(x + \sqrt{x^{2} + 1}) if − ∞ < x < ∞

— * —

)clear all
--S 166 of 298
asinhrule:=rule(asinh(x) == log(x+sqrt(x^2+1)))
--R
--R                         +------+
--R                         | 2
--R   (1)  asinh(x) == log(\|x  + 1  + x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 166
--S 167 of 298
t1:=asinh(x) - log(x+sqrt(x^2+1))
--R
--R               +------+
--R               | 2
--R   (2)  - log(\|x  + 1  + x) + asinh(x)
--R                                                    Type: Expression(Integer)
--E 167
--S 168 of 298
t2:=asinhrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 168

———-

10.2 8.56 cosh^{−1}(x)

acosh(x) == log(x + \sqrt{x^{2} − 1}) if x ≥ 1

)clear all
--S 169 of 298
acoshrule:=rule(acosh(x) == log(x+sqrt(x^2-1)))
--R
--R                         +------+
--R                         | 2
--R   (1)  acosh(x) == log(\|x  - 1  + x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 169
--S 170 of 298
t1:=acosh(x) - log(x+sqrt(x^2-1))
--R
--R               +------+
--R               | 2
--R   (2)  - log(\|x  - 1  + x) + acosh(x)
--R                                                    Type: Expression(Integer)
--E 170
--S 171 of 298
t2:=acoshrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 171

———-

10.3 8.57 tanh^{−1}(x)

atanh(x) == \frac{1}{2} log(\frac{1 + x}{1 − x}) if − 1 < x < 1

— * —

)clear all
--S 172 of 298
atanhrule:=rule(atanh(x) == 1/2*log((1+x)/(1-x)))
--R
--R                        - x - 1
--R                    log(-------)
--R                         x - 1
--R   (1)  atanh(x) == ------------

--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 172
--S 173 of 298
t1:=atanh(x) - 1/2*log((1+x)/(1-x))
--R
--R              - x - 1
--R        - log(-------) + 2atanh(x)
--R               x - 1
--R   (2)  --------------------------
--R                     2
--R                                                    Type: Expression(Integer)
--E 173
--S 174 of 298
t2:=atanhrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 174

———-

10.4 8.58 coth^{−1}(x)

acoth(x) == \frac{1}{2} log(\frac{x + 1}{x − 1}) if x > 1orx < −1

— * —

)clear all
--S 175 of 298
acothrule:=rule(acoth(x) == 1/2*log((x+1)/(x-1)))
--R
--R                        x + 1
--R                    log(-----)
--R                        x - 1
--R   (1)  acoth(x) == ----------
--R                         2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 175
--S 176 of 298
t1:=acoth(x) - 1/2*log((x+1)/(x-1))
--R
--R              x + 1

--R        - log(-----) + 2acoth(x)
--R              x - 1
--R   (2)  ------------------------
--R                    2
--R                                                    Type: Expression(Integer)
--E 176
--S 177 of 298
t2:=acothrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 177

———-

10.5 8.59 sech^{−1}(x)

asech(x) == log(\frac{1}{x} + \sqrt{\frac{1}{x^{2}} − 1}) if 0 < x ≤ 1

— * —

)clear all
--S 178 of 298
asechrule:=rule(asech(x) == log(1/x+sqrt(1/x^2-1)))
--R
--R                          +--------+
--R                          |   2
--R                          |- x  + 1
--R                        x |--------  + 1
--R                          |    2
--R                         \|   x
--R   (1)  asech(x) == log(----------------)
--R                                x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 178
--S 179 of 298
t1:=asech(x) - log(1/x+sqrt(1/x^2-1))
--R
--R                +--------+
--R                |   2
--R                |- x  + 1
--R              x |--------  + 1
--R                |    2
--R               \|   x

--R   (2)  - log(----------------) + asech(x)
--R                      x
--R                                                    Type: Expression(Integer)
--E 179
--S 180 of 298
t2:=asechrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 180

———-

10.6 8.60 csch^{−1}(x)

acsch(x) == log(\frac{1}{x} + \sqrt{\frac{1}{x^{2}} + 1}) if x ≠ 0

— * —

)clear all
--S 181 of 298
acschrule:=rule(acsch(x) == log(1/x+sqrt(1/x^2+1)))
--R
--R                          +------+
--R                          | 2
--R                          |x  + 1
--R                        x |------  + 1
--R                          |   2
--R                         \|  x
--R   (1)  acsch(x) == log(--------------)
--R                               x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 181
--S 182 of 298
t1:=acsch(x) - log(1/x+sqrt(1/x^2+1))
--R
--R                +------+
--R                | 2
--R                |x  + 1
--R              x |------  + 1
--R                |   2
--R               \|  x
--R   (2)  - log(--------------) + acsch(x)
--R                     x

--S 183 of 298
t2:=acschrule t1
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E 183

———-

11 Relations between Inverse Hyperbolic Functions

11.1 8.61 csch^{−1}(x)

csch^{−1}(x) == sinh^{−1} (\frac{1}{x})

— * —

)clear all
--S 184 of 298
cschinv:=rule(csch(x)^(-1) == sinh(1/x)^(-1))
--R
--R           1          1
--R   (1)  ------- == -------
--R        csch(x)         1
--R                   sinh(-)
--R                        x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 184
--S 185 of 298
t1:=csch(x)^(-1) - sinh(1/x)^(-1)
--R
--R             1
--R        sinh(-) - csch(x)
--R             x
--R   (2)  -----------------
--R                      1
--R          csch(x)sinh(-)
--R                      x
--R                                                    Type: Expression(Integer)
--E 185

--S 186 of 298
t2:=cschinv t1
--R
--R             1
--R        sinh(-) - csch(x)
--R             x
--R   (3)  -----------------
--R                      1
--R          csch(x)sinh(-)
--R                      x
--R                                                    Type: Expression(Integer)
--E 186

———-

11.2 8.62 sech^{−1}(x)

sech^{−1}(x) == cosh^{−1} (\frac{1}{x})

— * —

)clear all
--S 187 of 298
sechinv:=rule(sech(x)^(-1) == cosh(1/x)^(-1))
--R
--R           1          1
--R   (1)  ------- == -------
--R        sech(x)         1
--R                   cosh(-)
--R                        x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 187
--S 188 of 298
t1:=sech(x)^(-1) - cosh(1/x)^(-1)
--R
--R                         1
--R        - sech(x) + cosh(-)
--R                         x
--R   (2)  -------------------
--R                1
--R           cosh(-)sech(x)
--R                x
--R                                                    Type: Expression(Integer)
--E 188

--S 189 of 298
t2:=sechinv t1
--R
--R                         1
--R        - sech(x) + cosh(-)
--R                         x
--R   (3)  -------------------
--R                1
--R           cosh(-)sech(x)
--R                x
--R                                                    Type: Expression(Integer)
--E 189

———-

11.3 8.63 coth^{−1}(x)

coth^{−1}(x) == tanh^{−1} (\frac{1}{x})

— * —

)clear all
--S 190 of 298
cothinv:=rule(coth(x)^(-1) == tanh(1/x)^(-1))
--R
--R           1          1
--R   (1)  ------- == -------
--R        coth(x)         1
--R                   tanh(-)
--R                        x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 190
--S 191 of 298
t1:=coth(x)^(-1) - tanh(1/x)^(-1)
--R
--R             1
--R        tanh(-) - coth(x)
--R             x
--R   (2)  -----------------
--R                      1
--R          coth(x)tanh(-)
--R                      x
--R                                                    Type: Expression(Integer)
--E 191

--S 192 of 298
t2:=cothinv t1
--R
--R             1
--R        tanh(-) - coth(x)
--R             x
--R   (3)  -----------------
--R                      1
--R          coth(x)tanh(-)
--R                      x
--R                                                    Type: Expression(Integer)
--E 192

———-

11.4 8.64 sinh^{−1}(x)

sinh^{−1}(x) == − sinh^{−1}(x)

These identities are already known to Axiom — * —

)clear all
--S 193 of 298
t1:=sinh(-x)^(-1) - -sinh(x)^(-1)
--R
--R   (1)  0
--R                                                    Type: Expression(Integer)
--E 193

———-

11.5 8.65 tanh^{−1}(x)

tanh^{−1}(x) == − tanh^{−1}(x)

— * —

)clear all
--S 194 of 298
t1:=tanh(-x)^(-1) - -tanh(x)^(-1)
--R
--R   (1)  0

———-

11.6 8.66 coth^{−1}(x)

coth^{−1}(x) == − coth^{−1}(x)

— * —

)clear all
--S 195 of 298
t1:=coth(-x)^(-1) - -coth(x)^(-1)
--R
--R   (1)  0
--R                                                    Type: Expression(Integer)
--E 195

———-

11.7 8.67 csch^{−1}(x)

csch^{−1}(x) == − csch^{−1}(x)

— * —

)clear all
--S 196 of 298
t1:=csch(-x)^(-1) - -csch(x)^(-1)
--R
--R   (1)  0
--R                                                    Type: Expression(Integer)
--E 196

———-

12 Relationship between Hyperbolic and Trigonometric Functions

12.1 8.74 sin(ix)

sin(ix) == i sinh(x)

This match does not work.

— * —

)clear all
--S 197 of 298
sininv:=rule(sin(%i*x) == %i*sinh(x))
--R
--R   (1)  sin(%i x) == %i sinh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 197
--S 198 of 298
t1:=sin(x*%i) - %i*sinh(x)
--R
--R   (2)  - %i sinh(x) + sin(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 198
--S 199 of 298
t2:=sininv t1
--R
--R   (3)  - %i sinh(x) + sin(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 199

———-

12.2 8.75 cos(ix)

cos(ix) == cosh(x)

— * —

)clear all
--S 200 of 298
cosinv:=rule(cos(x*%i) == cosh(x))
--R

--R   (1)  cos(%i x) == cosh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 200
--S 201 of 298
t1:=cos(x*%i) - cosh(x)
--R
--R   (2)  - cosh(x) + cos(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 201
--S 202 of 298
t2:=cosinv t1
--R
--R   (3)  - cosh(x) + cos(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 202

———-

12.3 8.76 tan(ix)

tan(ix) == i tanh(x)

— * —

)clear all
--S 203 of 298
taninv:=rule(tan(x*%i) == %i*tanh(x))
--R
--R   (1)  tan(%i x) == %i tanh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 203
--S 204 of 298
t1:=tan(x*%i) - %i*tanh(x)
--R
--R   (2)  - %i tanh(x) + tan(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 204
--S 205 of 298
t2:=taninv t1
--R
--R   (3)  - %i tanh(x) + tan(%i x)
--R                                           Type: Expression(Complex(Integer))

--E 205

———-

12.4 8.77 csc(ix)

csc(ix) == −i csch(x)

— * —

)clear all
--S 206 of 298
cscinv:=rule(csc(x*%i) == -%i*csch(x))
--R
--R   (1)  csc(%i x) == - %i csch(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 206
--S 207 of 298
t1:=csc(x*%i) - -%i*csch(x)
--R
--R   (2)  %i csch(x) + csc(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 207
--S 208 of 298
t2:=cscinv t1
--R
--R   (3)  %i csch(x) + csc(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 208

———-

12.5 8.78 sec(ix)

sec(ix) == sech(x)

— * —

)clear all
--S 209 of 298

secinv:=rule(sec(x*%i) == sech(x))
--R
--R   (1)  sec(%i x) == sech(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 209
--S 210 of 298
t1:=sec(x*%i) - sech(x)
--R
--R   (2)  - sech(x) + sec(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 210
--S 211 of 298
t2:=secinv t1
--R
--R   (3)  - sech(x) + sec(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 211

———-

12.6 8.79 cot(ix)

cot(ix) == −i coth(x)

— * —

)clear all
--S 212 of 298
cotinv:=rule(cot(x*%i) == -%i*coth(x))
--R
--R   (1)  cot(%i x) == - %i coth(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 212
--S 213 of 298
t1:=cot(x*%i) - -%i*coth(x)
--R
--R   (2)  %i coth(x) + cot(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 213
--S 214 of 298
t2:=cotinv t1
--R

--R   (3)  %i coth(x) + cot(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 214

———-

12.7 8.80 sinh(ix)

sinh(ix) == i sin(x)

— * —

)clear all
--S 215 of 298
sinhinv:=rule(sinh(x*%i) == %i*sin(x))
--R
--R   (1)  sinh(%i x) == %i sin(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 215
--S 216 of 298
t1:=sinh(x*%i) - %i*sin(x)
--R
--R   (2)  sinh(%i x) - %i sin(x)
--R                                           Type: Expression(Complex(Integer))
--E 216
--S 217 of 298
t2:=sinhinv t1
--R
--R   (3)  sinh(%i x) - %i sin(x)
--R                                           Type: Expression(Complex(Integer))
--E 217

———-

12.8 8.81 cosh(ix)

cosh(ix) == cos(x)

— * —

--S 218 of 298
coshinv:=rule(cosh(x*%i) == cos(x))
--R
--R   (1)  cosh(%i x) == cos(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 218
--S 219 of 298
t1:=cosh(x*%i) - cos(x)
--R
--R   (2)  cosh(%i x) - cos(x)
--R                                           Type: Expression(Complex(Integer))
--E 219
--S 220 of 298
t2:=coshinv t1
--R
--R   (3)  cosh(%i x) - cos(x)
--R                                           Type: Expression(Complex(Integer))
--E 220

———-

p

12.9 8.82 tanh(ix)

tanh(ix) == i tan(x)

— * —

)clear all
--S 221 of 298
tanhinv:=rule(tanh(x*%i) == %i*tan(x))
--R
--R   (1)  tanh(%i x) == %i tan(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 221
--S 222 of 298
t1:=tanh(x*%i) - %i*tan(x)
--R
--R   (2)  tanh(%i x) - %i tan(x)
--R                                           Type: Expression(Complex(Integer))
--E 222

--S 223 of 298
t2:=tanhinv t1
--R
--R   (3)  tanh(%i x) - %i tan(x)
--R                                           Type: Expression(Complex(Integer))
--E 223

———-

12.10 8.83 csch(ix)

csch(ix) == −i csc(x)

— * —

)clear all
--S 224 of 298
cschinv:=rule(x*%i == -%i*csc(x))
--R
--R   (1)  %i x == - %i csc(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 224
--S 225 of 298
t1:=x*%i - -%i*csc(x)
--R
--R   (2)  %i csc(x) + %i x
--R                                           Type: Expression(Complex(Integer))
--E 225
--S 226 of 298
t2:=cschinv t1
--R
--R   (3)  %i csc(x) + %i x
--R                                           Type: Expression(Complex(Integer))
--E 226

———-

12.11 8.84 sech(ix)

sech(ix) == sec(x)

— * —

--S 227 of 298
sechinv:=rule(sech(x*%i) == sec(x))
--R
--R   (1)  sech(%i x) == sec(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 227
--S 228 of 298
t1:=sech(x*%i) - sec(x)
--R
--R   (2)  sech(%i x) - sec(x)
--R                                           Type: Expression(Complex(Integer))
--E 228
--S 229 of 298
t2:=sechinv t1
--R
--R   (3)  sech(%i x) - sec(x)
--R                                           Type: Expression(Complex(Integer))
--E 229

———-

12.12 8.85 coth(ix)

coth(ix) == −i cot(x)

— * —

)clear all
--S 230 of 298
cothinv:=rule(coth(x*%i) == -%i*cot(x))
--R
--R   (1)  coth(%i x) == - %i cot(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 230
--S 231 of 298
t1:=coth(x*%i) - -%i*cot(x)
--R
--R   (2)  coth(%i x) + %i cot(x)
--R                                           Type: Expression(Complex(Integer))
--E 231

--S 232 of 298
t2:=cothinv t1
--R
--R   (3)  coth(%i x) + %i cot(x)
--R                                           Type: Expression(Complex(Integer))
--E 232

———-

13 Periodicity of Hyperbolic Functions

13.1 8.86 sinh(x + 2kπi)

sinh(x + 2kπi) == sinh(x)

— * —

)clear all
--S 233 of 298
sinhperiod:=rule(sinh(x+2*k*%pi*%i) == sinh(x))
--R
--R   (1)  sinh(x + 2%i k %pi) == sinh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 233
--S 234 of 298
t1:=sinh(x+2*k*%pi*%i) - sinh(x)
--R
--R   (2)  sinh(x + 2%i k %pi) - sinh(x)
--R                                           Type: Expression(Complex(Integer))
--E 234
--S 235 of 298
t2:=sinhperiod t1
--R
--R   (3)  sinh(x + 2%i k %pi) - sinh(x)
--R                                           Type: Expression(Complex(Integer))
--E 235

———-

13.2 8.87 cosh(x + 2kπi)

cosh(x + 2kπi) == cosh(x)

)clear all
--S 236 of 298
coshperiod:=rule(cosh(x+2*k*%pi*%i) == cosh(x))
--R
--R   (1)  cosh(x + 2%i k %pi) == cosh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 236
--S 237 of 298
t1:=cosh(x+2*k*%pi*%i) - cosh(x)
--R
--R   (2)  cosh(x + 2%i k %pi) - cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 237
--S 238 of 298
t2:=coshperiod t1
--R
--R   (3)  cosh(x + 2%i k %pi) - cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 238

———-

13.3 8.88 tanh(x + 2kπi)

tanh(x + kπi) == tanh(x)

— * —

)clear all
--S 239 of 298
tanhperiod:=rule(tanh(x+k*%pi*%i) == tanh(x))
--R
--R   (1)  tanh(x + %i k %pi) == tanh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 239
--S 240 of 298
t1:=tanh(x+k*%pi*%i) - tanh(x)
--R
--R   (2)  tanh(x + %i k %pi) - tanh(x)

--R                                           Type: Expression(Complex(Integer))
--E 240
--S 241 of 298
t2:=tanhperiod t1
--R
--R   (3)  tanh(x + %i k %pi) - tanh(x)
--R                                           Type: Expression(Complex(Integer))
--E 241

———-

13.4 8.89 csch(x + 2kπi)

csch(x + 2kπi) == csch(x)

— * —

)clear all
--S 242 of 298
cschperiod:=rule(csch(x+2*k*%pi*%i) == csch(x))
--R
--R   (1)  csch(x + 2%i k %pi) == csch(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 242
--S 243 of 298
t1:=csch(x+2*k*%pi*%i) - csch(x)
--R
--R   (2)  csch(x + 2%i k %pi) - csch(x)
--R                                           Type: Expression(Complex(Integer))
--E 243
--S 244 of 298
t2:=cschperiod t1
--R
--R   (3)  csch(x + 2%i k %pi) - csch(x)
--R                                           Type: Expression(Complex(Integer))
--E 244

———-

13.5 8.90 sech(x + 2kπi)

sech(x + 2kπi) == sech(x)

)clear all
--S 245 of 298
sechperiod:=rule(sech(x+2*k*%pi*%i) == sech(x))
--R
--R   (1)  sech(x + 2%i k %pi) == sech(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 245
--S 246 of 298
t1:=sech(x+2*k*%pi*%i) - sech(x)
--R
--R   (2)  sech(x + 2%i k %pi) - sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 246
--S 247 of 298
t2:=sechperiod t1
--R
--R   (3)  sech(x + 2%i k %pi) - sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 247

———-

13.6 8.91 coth(x + 2kπi)

coth(x + kπi) == coth(x)

— * —

)clear all
--S 248 of 298
cothperiod:=rule(coth(x+k*%pi*%i) == coth(x))
--R
--R   (1)  coth(x + %i k %pi) == coth(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 248
--S 249 of 298
t1:=coth(x+k*%pi*%i) - coth(x)
--R
--R   (2)  coth(x + %i k %pi) - coth(x)

--S 250 of 298
t2:=cothperiod t1
--R
--R   (3)  coth(x + %i k %pi) - coth(x)
--R                                           Type: Expression(Complex(Integer))
--E 250

———-

14 Inverse Hyperbolic and Inverse Trig Functions

These patterns do not match in Axiom.

14.1 8.92 sin^{−1}(ix)

sin^{−1}(ix) == i sinh^{−1}(x)

— * —

)clear all
--S 251 of 298
sinsinh:=rule(sin(%i*x)^(-1) == %i*sinh(x)^(-1))
--R
--R            1           %i
--R   (1)  --------- == -------
--R        sin(%i x)    sinh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 251
--S 252 of 298
t1:=sin(%i*x)^(-1) - %i*sinh(x)^(-1)
--R
--R        sinh(x) - %i sin(%i x)
--R   (2)  ----------------------
--R           sin(%i x)sinh(x)
--R                                           Type: Expression(Complex(Integer))
--E 252
--S 253 of 298
t2:=sinsinh t1
--R
--R        sinh(x) - %i sin(%i x)

--R   (3)  ----------------------
--R           sin(%i x)sinh(x)
--R                                           Type: Expression(Complex(Integer))
--E 253

———-

14.2 8.93 sinh^{−1}(ix)

sinh^{−1}(ix) == i sin^{−1}(x)

— * —

)clear all
--S 254 of 298
sinhsin:=rule(sinh(%i*x)^(-1) == %i*sin(x)^(-1))
--R
--R             1          %i
--R   (1)  ---------- == ------
--R        sinh(%i x)    sin(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 254
--S 255 of 298
t1:=sinh(%i*x)^(-1) - %i*sin(x)^(-1)
--R
--R        - %i sinh(%i x) + sin(x)
--R   (2)  ------------------------
--R            sin(x)sinh(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 255
--S 256 of 298
t2:=sinhsin t1
--R
--R        - %i sinh(%i x) + sin(x)
--R   (3)  ------------------------
--R            sin(x)sinh(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 256

———-

14.3 8.94 cos^{−1}(x)

cos^{−1}(x) == ± cosh^{−1}(x)

)clear all
--S 257 of 298
coscosh:=rule(cos(x)^(-1) == %i*cosh(x)^(-1))
--R
--R           1         %i
--R   (1)  ------ == -------
--R        cos(x)    cosh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 257
--S 258 of 298
t1:=cos(x)^(-1) - %i*cosh(x)^(-1)
--R
--R        cosh(x) - %i cos(x)
--R   (2)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 258
--S 259 of 298
t2:=coscosh t1
--R
--R        cosh(x) - %i cos(x)
--R   (3)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 259
)clear all
--S 260 of 298
coscosh2:=rule(cos(x)^(-1) == -%i*cosh(x)^(-1))
--R
--R           1           %i
--R   (1)  ------ == - -------
--R        cos(x)      cosh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 260
--S 261 of 298
t1:=cos(x)^(-1) - -%i*cosh(x)^(-1)
--R
--R        cosh(x) + %i cos(x)
--R   (2)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))

--E 261
--S 262 of 298
t2:=coscosh2 t1
--R
--R        cosh(x) + %i cos(x)
--R   (3)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 262

———-

14.4 8.95 cosh^{−1}(x)

cosh^{−1}(x) == ±i cos^{−1}(x)

— * —

)clear all
--S 263 of 298
coshcos:=rule(cosh(x)^(-1) == %i*cos(x)^(-1))
--R
--R           1         %i
--R   (1)  ------- == ------
--R        cosh(x)    cos(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 263
--S 264 of 298
t1:=cosh(x)^(-1) - %i*cos(x)^(-1)
--R
--R        - %i cosh(x) + cos(x)
--R   (2)  ---------------------
--R            cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 264
--S 265 of 298
t2:=coshcos t1
--R
--R        - %i cosh(x) + cos(x)
--R   (3)  ---------------------
--R            cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 265

--S 266 of 298
coshcos2:=rule(cosh(x)^(-1) == -%i*cos(x)^(-1))
--R
--R           1           %i
--R   (1)  ------- == - ------
--R        cosh(x)      cos(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 266
--S 267 of 298
t1:=cosh(x)^(-1) - -%i*cos(x)^(-1)
--R
--R        %i cosh(x) + cos(x)
--R   (2)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 267
--S 268 of 298
t2:=coshcos2 t1
--R
--R        %i cosh(x) + cos(x)
--R   (3)  -------------------
--R           cos(x)cosh(x)
--R                                           Type: Expression(Complex(Integer))
--E 268

———-

14.5 8.96 tan^{−1}(ix)

tan^{−1}(ix) == i tanh^{−1}(x)

— * —

)clear all
--S 269 of 298
tantanh:=rule(tan(%i*x)^(-1) == %i*tanh(x)^(-1))
--R
--R            1           %i
--R   (1)  --------- == -------
--R        tan(%i x)    tanh(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))

--E 269
--S 270 of 298
t1:=tan(%i*x)^(-1) - %i*tanh(x)^(-1)
--R
--R        tanh(x) - %i tan(%i x)
--R   (2)  ----------------------
--R           tan(%i x)tanh(x)
--R                                           Type: Expression(Complex(Integer))
--E 270
--S 271 of 298
t2:=tantanh t1
--R
--R        tanh(x) - %i tan(%i x)
--R   (3)  ----------------------
--R           tan(%i x)tanh(x)
--R                                           Type: Expression(Complex(Integer))
--E 271

———-

14.6 8.97 tanh^{−1}(ix)

tanh^{−1}(ix) == i tan^{−1}(x)

— * —

)clear all
--S 272 of 298
tanhtan:=rule(tanh(%i*x)^(-1) == %i*tan(x)^(-1))
--R
--R             1          %i
--R   (1)  ---------- == ------
--R        tanh(%i x)    tan(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 272
--S 273 of 298
t1:=tanh(%i*x)^(-1) - %i*tan(x)^(-1)
--R
--R        - %i tanh(%i x) + tan(x)
--R   (2)  ------------------------
--R            tan(x)tanh(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 273

--S 274 of 298
t2:=tanhtan t1
--R
--R        - %i tanh(%i x) + tan(x)
--R   (3)  ------------------------
--R            tan(x)tanh(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 274

———-

14.7 8.98 cot^{−1}(ix)

cot^{−1}(ix) == −i cosh^{−1}(x)

— * —

)clear all
--S 275 of 298
cotcoth:=rule(cot(%i*x)^(-1) == -%i*coth(x)^(-1))
--R
--R            1             %i
--R   (1)  --------- == - -------
--R        cot(%i x)      coth(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 275
--S 276 of 298
t1:=cot(%i*x)^(-1) - -%i*coth(x)^(-1)
--R
--R        coth(x) + %i cot(%i x)
--R   (2)  ----------------------
--R           cot(%i x)coth(x)
--R                                           Type: Expression(Complex(Integer))
--E 276
--S 277 of 298
t2:=cotcoth t1
--R
--R        coth(x) + %i cot(%i x)
--R   (3)  ----------------------
--R           cot(%i x)coth(x)
--R                                           Type: Expression(Complex(Integer))
--E 277

———-

14.8 8.99 coth^{−1}(ix)

coth^{−1}(ix) == −i cot^{−1}(x)

— * —

)clear all
--S 278 of 298
cothcot:=rule(coth(%i*x)^(-1) == -%i*cot(x)^(-1))
--R
--R             1            %i
--R   (1)  ---------- == - ------
--R        coth(%i x)      cot(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 278
--S 279 of 298
t1:=coth(%i*x)^(-1) - -%i*cot(x)^(-1)
--R
--R        %i coth(%i x) + cot(x)
--R   (2)  ----------------------
--R           cot(x)coth(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 279
--S 280 of 298
t2:=cothcot t1
--R
--R        %i coth(%i x) + cot(x)
--R   (3)  ----------------------
--R           cot(x)coth(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 280

———-

14.9 8.100 sec^{−1}(x)

sec^{−1}(x) == ±i sech(x)

— * —

)clear all

--S 281 of 298
secsech:=rule(sec(x)^(-1) == %i*sech(x)^(-1))
--R
--R           1         %i
--R   (1)  ------ == -------
--R        sec(x)    sech(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 281
--S 282 of 298
t1:=sec(x)^(-1) - %i*sech(x)^(-1)
--R
--R        sech(x) - %i sec(x)
--R   (2)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 282
--S 283 of 298
t2:=secsech t1
--R
--R        sech(x) - %i sec(x)
--R   (3)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 283
)clear all
--S 284 of 298
secsech2:=rule(sec(x)^(-1) == -%i*sech(x)^(-1))
--R
--R           1           %i
--R   (1)  ------ == - -------
--R        sec(x)      sech(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 284
--S 285 of 298
t1:=sec(x)^(-1) - -%i*sech(x)^(-1)
--R
--R        sech(x) + %i sec(x)
--R   (2)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 285
--S 286 of 298
t2:=secsech2 t1
--R

--R        sech(x) + %i sec(x)
--R   (3)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 286

———-

14.10 8.101 sech^{−1}(x)

sech^{−1}(x) == ±i sec^{−1}(x)

— * —

)clear all
--S 287 of 298
sechsec:=rule(sech(x)^(-1) == %i*sec(x)^(-1))
--R
--R           1         %i
--R   (1)  ------- == ------
--R        sech(x)    sec(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 287
--S 288 of 298
t1:=sech(x)^(-1) - %i*sec(x)^(-1)
--R
--R        - %i sech(x) + sec(x)
--R   (2)  ---------------------
--R            sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 288
--S 289 of 298
t2:=sechsec t1
--R
--R        - %i sech(x) + sec(x)
--R   (3)  ---------------------
--R            sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 289
)clear all
--S 290 of 298
sechsec:=rule(sech(x)^(-1) == -%i*sec(x)^(-1))

--R
--R           1           %i
--R   (1)  ------- == - ------
--R        sech(x)      sec(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 290
--S 291 of 298
t1:=sech(x)^(-1) - -%i*sec(x)^(-1)
--R
--R        %i sech(x) + sec(x)
--R   (2)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 291
--S 292 of 298
t2:=sechsec t1
--R
--R        %i sech(x) + sec(x)
--R   (3)  -------------------
--R           sec(x)sech(x)
--R                                           Type: Expression(Complex(Integer))
--E 292

———-

14.11 8.102 csc^{−1}(ix)

csc^{−1}(ix) == −i csch^{−1}(x)

— * —

)clear all
--S 293 of 298
csccsch:=rule(csc(%i*x)^(-1) == -%i*csch(x)^(-1))
--R
--R            1             %i
--R   (1)  --------- == - -------
--R        csc(%i x)      csch(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 293
--S 294 of 298
t1:=csc(%i*x)^(-1) - -%i*csch(x)^(-1)
--R

--R        csch(x) + %i csc(%i x)
--R   (2)  ----------------------
--R           csc(%i x)csch(x)
--R                                           Type: Expression(Complex(Integer))
--E 294
--S 295 of 298
t2:=csccsch t1
--R
--R        csch(x) + %i csc(%i x)
--R   (3)  ----------------------
--R           csc(%i x)csch(x)
--R                                           Type: Expression(Complex(Integer))
--E 295

———-

14.12 8.103 csch^{−1}(ix)

csch^{−1}(ix) == −i csc^{−1}(x)

— * —

)clear all
--S 296 of 298
cschcsc:=rule(csch(%i*x)^(-1) == -%i*csc(x)^(-1))
--R
--R             1            %i
--R   (1)  ---------- == - ------
--R        csch(%i x)      csc(x)
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E 296
--S 297 of 298
t1:=csch(%i*x)^(-1) - -%i*csc(x)^(-1)
--R
--R        %i csch(%i x) + csc(x)
--R   (2)  ----------------------
--R           csc(x)csch(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 297
--S 298 of 298
t2:=cschcsc t1
--R
--R        %i csch(%i x) + csc(x)

--R   (3)  ----------------------
--R           csc(x)csch(%i x)
--R                                           Type: Expression(Complex(Integer))
--E 298

———-

rootrule4a:RewriteRule(INT,COMPLEX(INT),EXPR(COMPLEX(INT))):=
  rule(sqrt(p^2-q^2)==sqrt(p-q)*sqrt(q+p))  schaum17.input.pamphlet:
tanrule2:RewriteRule(INT,COMPLEX(INT),EXPR(COMPLEX(INT))):=
  rule(tan(a) == sin(a)/cos(a))  schaum17.input.pamphlet:

— * —

)spool
)lisp (bye)

———-

References^{[1] nothing}

79